# Stabilizing black-box model selection with the inflated argmax

### v1

@@ Appendix B Experiment details
In both our experiments, we utilize a cluster computing system to distribute parallel jobs across CPU nodes.
B.1 Regression example
This section provides further details about the experiment in § 5.1, including the data generation process and a justification for the regularization hyperparameter $\lambda$.
B.1.1 Data generation
$x_{i}\sim\mathcal{N}(0,C),$ (18)
where $x_{i}\in\mathbb{R}^{d}$, $\mathcal{N}$ denotes a Normal distribution, and $C\in\mathbb{R}^{d\times d}$ is visualized in Figure 1. The nonzero off-diagonal elements of $C$ are 0.99, corresponding to extremely correlated covariates. For a single trial $j$, we simulate a dataset with $d=20$ covariates and $i\in[n]$ ($n=30$) samples. We take the first and third columns as the covariates used to create the response $y$,
B.1.2 Optimization hyperparameter
B.2 Lotka-Volterra simulations
@@ B.2.1 Data generation
B.2.2 Cross-validating optimization hyperparameters
@@
We provide a table of validation MSEs averaged across the 5 folds in Table 2. Given our discretization of $\lambda$ and $\omega$ values, we choose $\omega=0.18$ and $\lambda=0.01$ and keep these values fixed in our experiments presented in § 5.1. Figure 6 visualizes the corresponding model selected by 5-fold cross validation. The best estimated model is
@@
Appendix C Stability with varying numbers of bags
Figure 7: Empirical CDF of instability measures $\delta_{j}$ across trials $j\in[N]$ in the regression experiment for varying number of bags $B$ across four different selection methods: the $\mathop{\!\arg\!\max}$, top-2, $\mathop{\!\arg\!\max}^{0.8}$, ip0.3. The values of $k$, $\varepsilon$, and $\tau$ were chosen to be consistent with the CDF plot in Figure 2(a).
Figure 8: Empirical CDF of instability measures $\delta_{j}$ across trials $j\in[N]$ in the Lotka-Volterra experiment for varying number of bags $B$ across four different selection methods: the argmax, top-2, argmax0.09, ip0.63. The values of $k$, $\varepsilon$, and $\tau$ were chosen to be consistent with the CDF plot in Figure 2(b).
The computational complexity of estimating the weights associated with each possible model $m\in M^{+}$ is not determined by the cardinality of the set $M^{+}$, but rather the number of bags $B$ used to estimated the weights associated with each model. Intuitively, we can generally expect that a larger number of bags (i.e., larger $B$) will correspond to a better estimate of the weights $\hat{w}$ for each model $m$ across selection methods. Figures 7 and 8 visualize the empirical CDF of instabilities $\delta_{j}$ across trials $j\in[N]$, where $N=200$ in the regression experiments and $N=100$ in the Lotka-Volterra experiment. A separate line is plotted for each choice of $B$.
@@ Appendix C Stability with varying numbers of bags
Appendix D Supplementary regression experiment
[…]
$\left\{\{x^{(1)},x^{(2)}\}\times\{x^{(3)},x^{(4)},x^{(5)}\}\right\}.$

### v2

@@ Appendix D Main experimental details
In our experiments, we utilize a cluster computing system to distribute parallel jobs across CPU nodes.
D.1 Lotka-Volterra simulations
@@ D.1.1 Data generation
D.1.2 Cross-validating optimization hyperparameters
@@
We provide a table of validation MSEs averaged across the 5 folds in Table 3. Given our discretization of $\lambda$ and $\omega$ values, we choose $\omega=0.18$ and $\lambda=0.01$ and keep these values fixed in our experiments presented in Appendix C. Figure 10 visualizes the corresponding model selected by 5-fold cross validation. The best estimated model is
@@
D.2 Graph subset selection
D.2.1 Cross-validation to select $\lambda$
We select the penalization hyperparameter $\lambda$ in (11) via 5-fold cross-validation. The validation log-likelihood averaged across the 5 folds is computed as
[…]
where $\hat{\Theta}_{v}$ is computed by solving (11) using the training data in fold index $v$, and $\tilde{\Sigma}_{v}$ is the sample covariance matrix computed from the validation data in fold index $v$. Computing (25) on the flow cytometry data across various choices of $\lambda\in[1,\dots,500]$ yields the curve shown in Figure 11. Based on this figure, the best choice of $\lambda$ is $\lambda=77$, which we keep constant throughout our experiments in §5.2.1.
Figure 11: Average 5-fold validation log-likelihood across different choices of $\lambda$. The $\lambda$ that maximizes the log-likelihood averaged over 5 held-out validation folds is $\lambda=77$.
Appendix E Stability with varying numbers of bags
Figure 12: Empirical CDF of instability measures $\delta_{j}$ across trials $j\in[N]$ in the regression experiment for varying number of bags $B$ across four different selection methods: the $\mathop{\!\arg\!\max}$, top-2, $\mathop{\!\arg\!\max}^{0.8}$, ip0.3. The values of $k$, $\varepsilon$, and $\tau$ were chosen to be consistent with the CDF plot in Figure 7.
Figure 13: Empirical CDF of instability measures $\delta_{j}$ across trials $j\in[N]$ in the Lotka-Volterra experiment for varying number of bags $B$ across four different selection methods: the argmax, top-2, argmax0.09, ip0.63. The values of $k$, $\varepsilon$, and $\tau$ were chosen to be consistent with the CDF plot in Figure 3.
The computational complexity of estimating the weights associated with each possible model $m\in M^{+}$ is not determined by the cardinality of the set $M^{+}$, but rather the number of bags $B$ used to estimated the weights associated with each model. Intuitively, we can generally expect that a larger number of bags (i.e., larger $B$) will correspond to a better estimate of the weights $\hat{w}$ for each model $m$ across selection methods. Figures 12 and 13 visualize the empirical CDF of instabilities $\delta_{j}$ across trials $j\in[N]$, where $N=200$ in the regression experiments and $N=100$ in the Lotka-Volterra experiment. A separate line is plotted for each choice of $B$.